OpenStudy (anonymous):

how do you do 8x^2 - 11x + 3 with the ax^2 +bx +c meathod

OpenStudy (badhi):

what do you mean by "do" is it to solve the equation, $$8x^2-11x+3=0$$

OpenStudy (anonymous):

yes i need help solving it

OpenStudy (anonymous):

The quadratic formula is \[x = \frac{ -b \pm \sqrt{b ^{2} - 4ac} }{ 2a }\] Where a, b, c are the coefficients of your equation In this case a = 8 b = -11 c = 3

OpenStudy (anonymous):

ok